OpenStudy (anonymous):

Keq question: PCL5(s) + H2O(g) -> 2HCl(g) + POCl3(g) Keq = 11 <- in a 4L volume container, the moles are: PCl5 = 0.012 mol H2O = 0.016mol HCl = 0.120 mol Calculate: [POCl3]

OpenStudy (anonymous):

My steps: Find Molarity for the known concentrations, so divide all the moles by 4. PCl5 = 1 M because it's a solid. Solids and Liquids have no concentration, so they are 1. H2o = 0.004 M HCl = 0.03^2 M Keq Equation = 0.03^2 [POCl3] ------------- = 11 0.004 x 1 I cannot find the proper fitting for the POCl3, every answer I get won't make the equation equal 11 when I plug in my numbers.

OpenStudy (abb0t):

I think you need to use an ICE-table for this. Are you familiar with that?